OpenStudy (anonymous):

lx-4l=lx-7l (Solve for x)

OpenStudy (anonymous):

step by step pls

OpenStudy (anonymous):

for solving equations like these, use graphs. |x-4| is a v-graph @ x=4, similarly |x-7| is a v-graph @ x=7, they intersect at x=11/2; or by normal method u hav, 2 cases x-4=x-7,which is an invalid opening. so the next opening is x-4=-x+7 which gives us , 2x=11 and hence x=11/2 got it ?

OpenStudy (anonymous):

yea. but howd u get -x+7?...u just multiplied negative by everything on that side?

OpenStudy (anonymous):

wen ever u open a |fn(x)| , u hav 2 cases, opening it as fn(x) or -fn(x) ;

OpenStudy (anonymous):

so here we hav x-7 or -(x-7)=-x+7

OpenStudy (anonymous):

so why didnt u do the same to lx-4l?

OpenStudy (anonymous):

you can try it and it will end with the same exact solution :)

OpenStudy (anonymous):

another way to do this is to square both sides x^2 -8x + 16 = x^2 -14x + 49 the x^2 disappear to leave a simple eqaution

OpenStudy (anonymous):

i actually did it, sorry i din't do it explicitly. since you hav 2 |fn(x)| u hav actually four combinations of openin, ++ +- -+ -- but case 1,4 and 2,3 are the same , so it looked as if i din't apply it on x-4, u cud've as well worked like -x+4=x-7 and end up getting the same answer :)

OpenStudy (anonymous):

ok. thanks